# CDI: Copyrighted Data Identification in Diffusion Models

### v2

@@ Abstract
Diffusion Models (DMs) benefit from large and diverse datasets for their training. Since this data is often scraped from the Internet without permission from the data owners, this raises concerns about copyright and intellectual property protections. While (illicit) use of data is easily detected for training samples perfectly re-created by a DM at inference time, it is much harder for data owners to verify if their data was used for training when the outputs from the suspect DM are not close replicas. Conceptually, membership inference attacks (MIAs), which detect if a given data point was used during training, present themselves as a suitable tool to address this challenge. However, we demonstrate that existing MIAs are not strong enough to reliably determine the membership of individual images in large, state-of-the-art DMs. To overcome this limitation, we propose Copyrighted Data Identification (CDI), a framework for data owners to identify whether their dataset was used to train a given DM. CDI relies on dataset inference techniques, i.e., instead of using the membership signal from a single data point, CDI leverages the fact that most data owners, such as providers of stock photography, visual media companies, or even individual artists, own datasets with multiple publicly exposed data points which might all be included in the training of a given DM. By selectively aggregating signals from existing MIAs and using new handcrafted methods to extract features from these datasets, feeding them to a scoring model, and applying rigorous statistical testing, CDI allows data owners with as little as 70 data points to identify with a confidence of more than 99% whether their data was used to train a given DM. Thereby, CDI represents a valuable tool for data owners to claim illegitimate use of their copyrighted data.
@@ 3 Limitations of MIAs in Member Detection
We rigorously evaluate existing MIAs to test their ability to detect training members in large, complex DMs. Prior studies [16, 26] reported success with MIAs in accurately identifying DMs training members; however, these results were often based on small-scale models or datasets (e.g., CIFAR100 [28]) that do not reflect the complexity of high-dimensional, diverse DM setups. Our analysis on state-of-the-art DMs trained on extensive datasets (i.e., ImageNet1k [14] or COCO [61]) reveals key performance limitations of existing MIAs and factors that contributed to overestimated effectiveness in previous work. Full details are provided in Appendix E.
@@ 3.2 Experimental Setup
We evaluate class-conditioned, as well as text-conditioned state-of-the-art DMs of various architectures, namely LDM [44], U-ViT [6], and DiT [36]. We employ already trained checkpoints provided by the respective papers [1, 4, 2]. For class-conditioned generative tasks, LDM offers one model checkpoint with the resolution of 256x256 (LDM256). For U-ViT and DiT, we have access to models operating on resolutions of 256x256 and 512x512 (U-ViT256, U-ViT512, DiT256, DiT512). Additionally, we conduct experiments on text-conditioned models based on U-ViT architecture (U-ViT256-T2I, U-ViT256-T2I-Deep) and a newly trained unconditional U-ViT256-Uncond model (see App. H).
@@ 3.3 Performance of MIAs in Member Detection
Our results indicate that the existing MIAs achieve performance comparable to random guessing. We present the aggregated max and average TPR@FPR=1% across 8 DMs in Tab. 1, and defer the full evaluation to App. R. For completeness, we provide AUC (Table 10), accuracy (Table 9), and ROC curves (Fig. 17) of MIAs there.
@@ 4 Our CDI Method
We design CDI as a tool for use in legal proceedings. Consequently, the CDI procedure is carried out by a third trusted party, referred to as an arbitrator. The arbitrator is approached by a victim, whose private data might have been potentially used in a training of a DM. The arbitrator executes CDI either in the gray-box model access, (can only obtain outputs, i.e., noise predictions, for given inputs to a DM at an arbitrary timestep $t$) or in the white-box model access (where DM’s internals and parameters can be inspected). The access type depends on the requirements of the features used in CDI (we provide more details in App. I).
@@ 4.1 Features
We provide further analysis of our features in App. Q.
@@ 4.3 Statistical Testing
Our choice of such a low $\alpha$ parameter is motivated by the TPR@FPR=1% metric established for MIAs [10], based on the intuition that the false positives are more harmful than false negatives in real-world applications, e.g., court cases. To improve the soundness of our statistical tests, we perform CDI 1000 times on randomly sampled subsets of P and U, and aggregate the obtained p-values [27, 62] (we provide more details in App. D).
@@ 5 Empirical Evaluation
Figure 3: Impact of feature selection. The values in the cells indicate the minimum size of P needed to reject $H_{0}$. Blue vertical lines separate results by the complexity of the feature set used to fit $s$: (left) our novel individual features from Sec. 4.1 and a joint existing MIAs feature, (second from left) all possible combinations of two, and of three features (second from right), and (right) the set of all available features.
@@ 5.1 Our CDI Confidently Identifies Collection of Data Samples as Training Data
We summarize the success of our CDI for diverse DMs and datasets in Fig. 2 (following the standard evaluation of DI as proposed in [35]). We report p-values as the confidence in the correct verification for different sizes of suspect data sets P. Our results highlight that CDI already enables a confident ($p<0.01$) dataset identification with as little as 70 samples (for instance, for U-ViT256-T2I-Deep DM trained on the COCO dataset) provided by the data owner. For DMs trained on larger datasets like ImageNet, we observe the need to increase the size of P to confidently reject the null hypothesis. More details on the impact of the size of P on the confidence of CDI can be found in App. F. In general, in our results, we identify the following trends: (1) For a given DM architecture, trained on given dataset, the number of samples required for the confident identification of the training data decreases with increasing input resolution (see App. J.2). (2) The larger the overall training set of the model, the more samples are needed for a confident claim (see also App. J.2). (3) The higher the number of model training steps the stronger the signal for identifying training data as shown in Fig. 6 in App. J.1.
@@ 5.2 Analysis of the Success of CDI
We perform multiple ablations on CDI’s building blocks to deepen understanding of its success. First, we show the importance of statistical testing as a core component of CDI. Then, we show that our features indeed boost the performance of CDI. Next, we demonstrate that our method remains effective even when not all samples from the suspect set were used in the DM’s training set. Finally, we show that CDI does not return false positives. We include additional evaluation of CDI in App. O and P, analysis of scoring model $s$ in App. M, and time complexity in App. H.2.
@@
A direct comparison between CDI’s p-values and TPR@FPR=1% for set-level MIA is challenging due to the differing metrics. To align them, we compute the power of the t-test with $\alpha=0.01$, which is equivalent to TPR@FPR=1% (see App. N). This approach allows us to directly compare set-level MIA to CDI without altering its methodology. The results in Tab. 2 underscore the critical role of statistical testing in CDI. Set-level MIA without statistical testing underperforms, while CDI with its rigorous testing achieves near-perfect performance for most of the models.
@@
We investigate CDI ’s behavior in cases where only a part of the samples in the suspect set P was used to train the DM, i.e., P contains a certain ratio of non-members, while the remaining samples in P are members. Practically, this corresponds to the situation where a data owner has a set of publicly exposed data points and suspects that all of them might have been used to train the DM, whereas, in reality, some were not used. This can happen, e.g., due to internal data cleaning on the side of the party who trained the DM. In particular, the data owner does not know which of their samples and how many of them have not been included into training. In Fig. 4 (extended by Fig. 7 in App. K), we present the success of CDI under different ratios of non-member samples in P. Note that our evaluation is the result of 1000 randomized experiments for each non-members ratio, model, and P size. We observe that CDI remains effective when the non-member ratio is over 0.5 and 0.8 for some models, i.e., it still correctly identifies P as training data of the model. Overall CDI’s robustness is higher when the data owner provides larger suspect datasets P. This is reflected in the p-value at the same non-member ratio decreasing as the number of samples in P increases.
@@
We analyze the effectiveness of performing CDI in the gray-box model access scenario, as defined in the threat model (Sec. 4). Therefore, we include only the original MIA features and Multiple Loss in CDI. Even in this case, CDI remains effective under gray-box model access and can reject the null hypothesis. In this more difficult scenario, across the eight tested DMs, CDI requires on average one-third more samples in P compared to the white-box access (where all features can be used). We refer to App. L for a detailed comparison.
@@ Acknowledgments
This work was supported by the German Research Foundation (DFG) within the framework of the Weave Programme under the project titled ”Protecting Creativity: On the Way to Safe Generative Models” with number 545047250. This research was also supported by the Polish National Science Centre (NCN) grant no. 2023/51/I/ST6/02854 and by the Warsaw University of Technology within the Excellence Initiative Research University (IDUB) programme. Responsibility for the content of this publication lies with the authors.
@@ Appendix A Broader Impact
Our research addresses the pressing issue of verifying the ownership of data used for training large image generation models, as highlighted by recent legal disputes [5]. By introducing CDI, we aim to enable data owners to verify if their data was (illegitimately) used for training DMs. Our work contributes to a more transparent and accountable ML ecosystem, aligning with broader societal values of fairness and respect for data ownership. We anticipate CDI will have a positive impact on both the ML community and society, promoting responsible and fair development of ML models.
@@ Appendix B Limitations
In this paper, we assess the effectiveness of CDI for image generation DMs. Although we believe that our methodology extends well to other modalities such as text or video, we do not perform an experimental evaluation in these areas. Our research focuses on diffusion models as the current state-of-the-art image generators widely employed in commercial applications. While acknowledging the existence of alternative generative image models, which have recently been shown to demonstrate comparable capabilities [24, 49], we choose to focus on DMs within the scope of our research. We note that CDI requires at least gray-box access, i.e. DM’s predictions at an arbitrary timestep $t$ to be effective. This limitation stems from the lack of reliable, strictly black-box MIA methods for DMs, i.e., methods that leverage only the final generated image, while avoiding the pitfalls described in App. E.
@@ C.1 Previous Membership Inference Attacks against DMs
An important consideration in PIA is that $f_{\theta}(z,0)$ follows a Gaussian, i.e., $f_{\theta}(z,0)\sim\mathcal{N}(0,I)$, as it should make the attack more performant. To assure that, Kong et al. [26] propose PIAN (normalized PIA), as an extension of their method. Formally, $\hat{f_{\theta}}(z,0)=C\cdot H\cdot W\sqrt{\frac{\pi}{2}}\frac{f_{\theta}(z,0)}{\lVert f_{\theta}(z,0)\rVert_{1}}$ with $C,H,W$ denoting the channels, height, and width of the input sample, respectively. However, our experiments in App. R and Sec. 5.1 indicate that this intuition does not translate to latent DMs, which is in line with findings from the original paper [26].
@@
Appendix D On Reporting p-values
@@
Appendix E On the Mismatch in MIAs Results
Contrary to the promising results of SecMIstat, PIA, and PIAN, our evaluation in App. R shows that these attacks fail to reach performance significantly higher than random guessing. The following is a methodological analysis of issues in the experimental setup proposed in [26, 16, 20]. We identify three pitfalls in their experimental settings: (1) usage of toy models, (2) overfitting to the evaluation set, and (3) distribution mismatch between members and non-members sets.
E.1 Pitfall 1: Toy Models
MIAs performance is directly correlated with the level of overfitting in the attacked model [52, 48]. Unfortunately, it is a common approach to evaluate the MIAs against DMs on very small toy models, trained on small-scale datasets like CIFAR10 [29], as it speeds up the inference. This in turn can lead to elevated performance which is further reported in published works [16, 26, 20], e.g., the TPR@FPR=1% at the level of 10% for SecMIstat and 30% for PIAN. We argue that evaluating any type of MIA in such setting is flawed, and provides incorrect insight on the performance of the proposed MIA.
E.2 Pitfall 2: Overfitting to the Evaluation Set
E.3 Pitfall 3: Mismatch in Distribution of Member and Non-Member Sets

### v3

@@ Abstract
Diffusion Models (DMs) benefit from large and diverse datasets for their training. Since this data is often scraped from the Internet without permission from the data owners, this raises concerns about copyright and intellectual property protections. While (illicit) use of data is easily detected for training samples perfectly re-created by a DM at inference time, it is much harder for data owners to verify if their data was used for training when the outputs from the suspect DM are not close replicas. Conceptually, membership inference attacks (MIAs), which detect if a given data point was used during training, present themselves as a suitable tool to address this challenge. However, we demonstrate that existing MIAs are not strong enough to reliably determine the membership of individual images in large, state-of-the-art DMs. To overcome this limitation, we propose Copyrighted Data Identification (CDI), a framework for data owners to identify whether their dataset was used to train a given DM. CDI relies on dataset inference techniques, i.e., instead of using the membership signal from a single data point, CDI leverages the fact that most data owners, such as providers of stock photography, visual media companies, or even individual artists, own datasets with multiple publicly exposed data points which might all be included in the training of a given DM. By selectively aggregating signals from existing MIAs and using new handcrafted methods to extract features from these datasets, feeding them to a scoring model, and applying rigorous statistical testing, CDI allows data owners with as little as 70 data points to identify with a confidence of more than 99% whether their data was used to train a given DM. Thereby, CDI represents a valuable tool for data owners to claim illegitimate use of their copyrighted data. We make our code available at https://github.com/sprintml/copyrighted_data_identification.
@@ 3 Limitations of MIAs in Member Detection
We rigorously evaluate existing MIAs to test their ability to detect training members in large, complex DMs. Prior studies [16, 27] reported success with MIAs in accurately identifying DMs training members; however, these results were often based on toy models or datasets (e.g., CIFAR100 [29]) that do not reflect the complexity of high-dimensional, diverse DM setups. Our analysis on state-of-the-art DMs trained on extensive datasets (i.e., ImageNet-1k [14] or COCO [64]) reveals significant performance limitations of existing MIAs and key factors that contributed to overestimated effectiveness in previous work. Full details are provided in Appendix D.
@@ 3.2 Experimental Setup
We evaluate class-conditioned, as well as text-conditioned state-of-the-art DMs of various architectures, namely LDM [47], U-ViT [6], and DiT [39]. We employ already trained checkpoints provided by the respective papers [1, 4, 2]. For class-conditioned generative tasks, LDM offers one model checkpoint with the resolution of 256x256 (LDM256). For U-ViT and DiT, we have access to models operating on resolutions of 256x256 and 512x512 (U-ViT256, U-ViT512, DiT256, DiT512). Additionally, we conduct experiments on text-conditioned models based on U-ViT architecture (U-ViT256-T2I, U-ViT256-T2I-Deep) and a newly trained unconditional U-ViT256-Uncond model (see App. I).
@@ 3.3 Performance of MIAs in Member Detection
Our results indicate that the existing MIAs achieve performance comparable to random guessing. We present the aggregated max and average TPR@FPR=1% across 8 DMs in Tab. 1, and defer the full evaluation to App. U. For completeness, we provide AUC (Table 14), accuracy (Table 13), and ROC curves (Fig. 17) of MIAs there.
@@ 4 Our CDI Method
We design CDI as a tool for use in legal proceedings. Consequently, the CDI procedure is carried out by a third trusted party, referred to as an arbitrator. The arbitrator is approached by a victim, whose private data might have been potentially used in a training of a DM. The arbitrator executes CDI either in the gray-box model access, (can only obtain outputs, i.e., noise predictions, for given inputs to a DM at an arbitrary timestep $t$) or in the white-box model access (where DM’s internals and parameters can be inspected). The access type depends on the requirements of the features used in CDI (we provide more details in App. N).
@@ 4.1 Features
We provide further analysis of our features in App. T.
@@ 4.3 Statistical Testing
Our choice of such a low $\alpha$ parameter is motivated by the TPR@FPR=1% metric established for MIAs [10], based on the intuition that the false positives are more harmful than false negatives in real-world applications, e.g., court cases. To improve the soundness of our statistical tests, we perform CDI 1000 times on randomly sampled subsets of P and U, and aggregate the obtained p-values [28, 65] (we provide more details in App. E).
@@ 5 Empirical Evaluation
Figure 3: Impact of feature selection. The values in the cells indicate the minimum size of P needed to reject $H_{0}$. Blue vertical lines separate results by the complexity of the feature set used to fit $s$: (left) our novel individual features from Sec. 4.1 and a joint existing MIAs feature, (second from left) all possible combinations of two, and of three features (second from right), and (right) all available features.
@@ 5.1 Our CDI Confidently Identifies Collection of Data Samples as Training Data
We summarize the success of our CDI for diverse DMs and datasets in Fig. 2 (following the standard evaluation of DI as proposed in [38]). We report p-values as the confidence in the correct verification for different sizes of suspect data sets P. Our results highlight that CDI already enables a confident ($p<0.01$) dataset identification with as little as 70 samples (for instance, for U-ViT256-T2I-Deep DM trained on the COCO dataset) provided by the data owner. For DMs trained on larger datasets like ImageNet, we observe the need to increase the size of P to confidently reject the null hypothesis. More details on the impact of the size of P on the confidence of CDI can be found in App. F. In general, in our results, we identify the following trends: (1) For a given DM architecture, trained on given dataset, the number of samples required for the confident identification of the training data decreases with increasing input resolution (see App. J). (2) The larger the overall training set of the model, the more samples are needed for a confident claim (see also App. J). (3) The higher the number of model training steps the stronger the signal for identifying training data as shown in Fig. 6 in App. K.
@@ 5.2 Analysis of the Success of CDI
We perform multiple ablations on CDI’s building blocks to deepen understanding of its success. First, we show the importance of statistical testing as a core component of CDI. Then, we show that our features indeed boost the performance of CDI. Next, we demonstrate that our method remains effective even when not all samples from the suspect set were used in the DM’s training set. Finally, we show that CDI does not return false positives. We include additional evaluation of CDI in App. R and S, analysis of scoring model $s$ in App. P, and time complexity in App. I.2. In Appendix H we evaluate CDI on additional DMs and showcase how CDI can be extended with additional feature extraction methods.
@@
A direct comparison between CDI’s p-values and TPR@FPR=1% for set-level MIA is challenging due to the differing metrics. To align them, we compute the power of the t-test with $\alpha=0.01$, which is equivalent to TPR@FPR=1% (see App. Q). This approach allows us to directly compare set-level MIA to CDI without altering its methodology. The results in Tab. 2 underscore the critical role of statistical testing in CDI. Set-level MIA without statistical testing underperforms, while CDI with its rigorous testing achieves near-perfect performance for most of the models.
@@
We investigate CDI ’s behavior in cases where only a part of the samples in the suspect set P was used to train the DM, i.e., P contains a certain ratio of non-members, while the remaining samples in P are members. Practically, this corresponds to the situation where a data owner has a set of publicly exposed data points and suspects that all of them might have been used to train the DM, whereas, in reality, some were not used. This can happen, e.g., due to internal data cleaning on the side of the party who trained the DM. In particular, the data owner does not know which of their samples and how many of them have not been included into training. In Fig. 4 (extended by Fig. 7 in App. L), we present the success of CDI under different ratios of non-member samples in P. Note that our evaluation is the result of 1000 randomized experiments for each non-members ratio, model, and P size. We observe that CDI remains effective when the non-member ratio is over 0.5 and 0.8 for some models, i.e., it still correctly identifies P as training data of the model. Overall CDI’s robustness is higher when the data owner provides larger suspect datasets P. This is reflected in the p-value at the same non-member ratio decreasing as the number of samples in P increases.
@@
We analyze the effectiveness of performing CDI in the gray-box model access scenario, as defined in the threat model (Sec. 4). Therefore, we include only the original MIA features and Multiple Loss in CDI. Even in this case, CDI remains effective under gray-box model access and can reject the null hypothesis. In this more difficult scenario, across the eight tested DMs, CDI requires on average one-third more samples in P compared to the white-box access (where all features can be used). We refer to App. O for a detailed comparison.
@@ Acknowledgments
This work was supported by the German Research Foundation (DFG) within the framework of the Weave Programme under the project titled ”Protecting Creativity: On the Way to Safe Generative Models” with number 545047250. This research was also supported by the Polish National Science Centre (NCN) grant no. 2023/51/I/ST6/02854 and 2020/39/O/ST6/01478 and by the Warsaw University of Technology within the Excellence Initiative Research University (IDUB) programme. Responsibility for the content of this publication lies with the authors.
@@ Appendix A Broader Impact
Our research addresses the pressing issue of verifying the ownership of data used for training large image generation models, as highlighted by recent legal disputes [5]. By introducing CDI, we aim to enable data owners to verify if their data was (illegitimately) used for training DMs. Our work contributes to a more transparent and accountable ML ecosystem, aligning with broader societal values of fairness and respect for data ownership. We anticipate CDI will have a positive impact on both the ML community and society, promoting responsible and fair development of ML models. We make our code available at https://github.com/sprintml/copyrighted_data_identification.
@@ Appendix B Limitations
In this paper, we assess the effectiveness of CDI for image generation DMs. Although we believe that our methodology extends well to other modalities such as text or video, we do not perform an experimental evaluation in these areas. Our research focuses on diffusion models as the current state-of-the-art image generators widely employed in commercial applications. While acknowledging the existence of alternative generative image models, which have recently been shown to demonstrate comparable capabilities [25, 52], we choose to focus on DMs within the scope of our research. We note that CDI requires at least gray-box access, i.e. DM’s predictions at an arbitrary timestep $t$ to be effective. This limitation stems from the lack of reliable, strictly black-box MIA methods for DMs, i.e., methods that leverage only the final generated image, while avoiding the pitfalls described in App. D.
@@ C.1 Previous Membership Inference Attacks against DMs
An important consideration in PIA is that $f_{\theta}(z,0)$ follows a Gaussian, i.e., $f_{\theta}(z,0)\sim\mathcal{N}(0,I)$, as it should make the attack more performant. To assure that, Kong et al. [27] propose PIAN (normalized PIA), as an extension of their method. Formally, $\hat{f_{\theta}}(z,0)=C\cdot H\cdot W\sqrt{\frac{\pi}{2}}\frac{f_{\theta}(z,0)}{\lVert f_{\theta}(z,0)\rVert_{1}}$ with $C,H,W$ denoting the channels, height, and width of the input sample, respectively. However, our experiments in App. U and Sec. 5.1 indicate that this intuition does not translate to latent DMs, which is in line with findings from the original paper [27].
@@
Appendix D On the Mismatch in MIAs Results
Contrary to the promising results of SecMIstat, PIA, and PIAN, our evaluation in App. U shows that these attacks fail to reach performance significantly higher than random guessing. The following is a methodological analysis of issues in the experimental setup proposed in [27, 16, 20]. We identify three pitfalls in their experimental settings: (1) usage of toy models, (2) overfitting to the evaluation set, and (3) distribution mismatch between members and non-members sets.
D.1 Pitfall 1: Toy Models
MIAs performance is directly correlated with the level of overfitting in the attacked model [55, 51]. Unfortunately, it is a common approach to evaluate the MIAs against DMs on very small toy models, trained or fine-tuned on small-scale datasets like CIFAR10 [30], as it speeds up the inference. This in turn can lead to elevated performance which is further reported in published works [16, 27, 20], e.g., the TPR@FPR=1% at the level of 10% for SecMIstat and 30% for PIAN. We argue that evaluating any type of MIA in such setting is flawed, and provides incorrect insight on the performance of the proposed MIA.
D.2 Pitfall 2: Overfitting to the Evaluation Set
D.3 Pitfall 3: Mismatch in Distribution of Member and Non-Member Sets
Appendix E On Reporting p-values
@@ Appendix G Additional Features
Appendix H Extending CDI
CDI provides an inherently flexible framework for identifying data collections used in the training of DMs. In this section, we discuss how CDI can be extended with additional feature extraction methods and applied to a broader range of DM architectures and data modalities.
H.1 Extending CDI with features from novel MIAs
As the field advances, CDI can be extended with additional feature extraction methods based on novel MIAs. We demonstrate this by incorporating CLiD [72] as a feature extraction method within CDI. First, in Table 4, we show that the performance of CLiD MIA remains limited for models trained on ImageNet (first five columns). While it significantly outperforms previous MIAs, as compared to Tab. 12, it still does not allow reliable membership identification for ImageNet-trained models. Note that we do not evaluate CLiD on U-ViT256-Uncond, as it requires both conditional and unconditional inputs to the DM.
Table 4: CLiD [72] MIA results at a TPR@FPR=1%. Values in the table are in %.
Model LDM U-ViT256 DiT-256 U-ViT512 DiT512 U-ViT256-T2I U-ViT256-T2I-Deep
TPR@FPR=1% 1.52 2.18 2.74 2.03 2.00 10.91 24.11
Next, we employ features from CLiD in CDI and present in Table 5 that it further improves our method. This confirms our premise from Section 4.2, that more meaningful features improve our method. It also demonstrates how CDI benefits from advancements in MIAs. When we extend CDI with CLiD, the method needs as few as 30 samples to identify data collections used in DM training.
Table 5: Effect of extending CDI with features extracted using CLiD. We report the minimal sample size of P needed to confidently reject the null hypothesis with CDI.
Model LDM U-ViT256 DiT256 U-ViT512 DiT512 U-ViT256-T2I U-ViT256-T2I-Deep
Base CDI 6000 3000 2000 2000 1000 200 70
With CLiD 4000 700 400 2000 700 50 30
H.2 Extending CDI to novel DMs
With new DM architectures introduced, CDI can be applied to a broader range of DMs. In Table 6, we evaluate the performance of CDI on SiT [37], MDT [21], and DiMR [34], and show that it successfully identifies the data collections used in DM training.
Table 6: Performance of CDI on additional DMs. We report the minimal sample size of P needed to confidently reject the null hypothesis with CDI.
Model SiT-XL/2 MDTv2-XL/2 MDTv1-XL/2 DiMR-XL/2R DiMR-G/2R
[…]
H.3 Extending CDI to other data modalities
CDI framework is inherently flexible and can be applied to data modalities beyond images, such as text or audio. This extension is straightforward due to the use of feature extraction methods that operate in the latent space of DMs.